OpenStudy (buttermequeasy):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (jamesj):

For example the equation x + 1 = 2. x is a solution of the equation if it satisfies the equation. Thus in this case x = 1 is a solution as indeed 1 + 1 = 2.

OpenStudy (anonymous):

if ax+bx=c then the number x=c/(a+b) is a solution and if you plug it in on the right hand side you should get c